looking to solve this question, I need different number combinations whose sum is equal to 133.
I need 7 numbers in each so...., 7 numbers between 1-47 whose sum is 133... I'm looking for all possible combinations...there should only be a couple..not sure how to go about doing it.

Answer by Edwin McCravy(20060)

There are too many unanswered questions about
what you are looking for.

Are you talking about 

sum 1+2+3+4+30+46+47 = 133 

all the way to sum:

16+17+18+19+20+21+22 = 133?

where there are 7 different numbers arranged from 
smallest to largest?

Do you allow repetitions such as

7+21+21+21+21+21+21 = 133?

Do you consider different permutations of the same 7
numbers as counted as separate? For instance do you consider
this:

1+2+3+4+30+46+47 = 133

different from

2+1+3+4+30+46+47 = 133

where only the 2 and the 1 are swapped? Or are they 
considered the same?

If you can let me know exactly what you want in the thank you 
note form below, I'll get back to you by email.  No charge
ever.

Edwin